OpenStudy (anonymous):

I'm back with more conversions, only this time pertaining to cost. You consult your inventory book and find the following information given as below: Dry Storage: Flour, All Purpose $30.00/bag 50#/bag Flour, Bread $32.50/bag 50#/bag Salt $1.14/container 12 oz/container Oil, Canola $54.00/cs 2 ea 17 1/2 # container/cs Cold Storage: Butter $2.13/# Eggs, Whole $23.20/tub 10#/tub Eggs, Yolks $27.80/tub 10#/tub Milk, Whole $3.75/ gal Using the information given (c)

OpenStudy (anonymous):

what does it cost to produce one recipe of egg noodles? (7# 11 oz)

OpenStudy (anonymous):

The previous recipe I was given was Egg Noodles: Bread Flour - 5# Eggs, Whole - 1# 8oz Salt - 2 oz Oil - 1 oz Water - 1 #

OpenStudy (anonymous):

I dont understand the "#". ??

OpenStudy (anonymous):

# = pounds

OpenStudy (anonymous):

Okay, thanks. Alright, hold on.

OpenStudy (anonymous):

No problem, thank you for your time.

OpenStudy (anonymous):

@Nnesha. lol I'm having a brain-fart.

OpenStudy (anonymous):

It's alright, there's a lot of information here.

OpenStudy (anonymous):

Sorry, I'm taking so long. lol its very tedious. XD Just a few more minutes.

jimthompson5910 (jim_thompson5910):

hint: try to connect the dollar value to the number of pounds (you'll have to take 2 steps)

OpenStudy (anonymous):

Do the pounds need to be converter to oz first? That's how they were with converting the recipes but it may be different here with cost...

jimthompson5910 (jim_thompson5910):

no you don't need it to be in pounds as long as the cost is in ounces

OpenStudy (anonymous):

Ah, yes! Correct. I didn't even think of that.

OpenStudy (anonymous):

Also, this goes without being said probably but I'm sure we only need to focus on the bread flour, salt, canola oil and whole eggs since they're the only things used within the recipe. At least that's what I'm assuming we focus on.

jimthompson5910 (jim_thompson5910):

Example: `Salt - 2 oz` `Salt $1.14/container 12 oz/container` we need 2 oz of salt there are `12 oz/container`; put another way 1 container = 12 oz so \[\Large (2 \ \text{oz})*\frac{1 \ \text{container}}{12 \ \text{oz}} = \frac{1}{6} \ \text{container}\] so saying "we need 2 ounces of salt" is the same as saying "we need 1/6 of a container of salt"

jimthompson5910 (jim_thompson5910):

`Salt $1.14/container` then we convert from "container" units to "dollars" units \[\Large \frac{1}{6} \ \text{container} * \frac{1.14 \ \text{dollars}}{1 \ \text{container}} = 0.19 \ \text{dollars}\] Summary: if we need 2 oz of salt, then that will cost $0.19 Now that I think about it, you probably can't buy 1/6 of a container, so you'll need to round that 1/6 to 1 and then convert "container" units to "dollars" units

OpenStudy (anonymous):

So, Bread Flour Bread Flour = 80 oz Bread Flour = $32.50/bag 50#/bag (80 oz) * 1 bag/800 oz = 1/10 bag?

jimthompson5910 (jim_thompson5910):

yes 80 oz = 1/10 of a bag but you probably can't buy a tenth of a bag. You'll have to buy the whole bag most likely

OpenStudy (anonymous):

1/10 bag * $32.50/1 bag = $3.25 ??

OpenStudy (anonymous):

I think it's going buy how much the ingredients would cost individually, otherwise we'd just buy one of everything and that'd be a little too obvious in my opinion.

jimthompson5910 (jim_thompson5910):

That's true. I was just trying to think of how it would apply in real life, but then again this hypothetical supermarket may make an exception to buy portions of bags, containers, etc

OpenStudy (anonymous):

This is a college cooking class and we had maybe 3 minutes of review for this, so I'd assume he'd want us to try and do some math even though it wasn't very thoroughly explained. x-x Still, do you think #3.25 for the bread flour was correct?

OpenStudy (anonymous):

*$3.25

jimthompson5910 (jim_thompson5910):

1/10 = 0.1 \[\Large \frac{0.1 \ \text{bag}}{1} * \frac{32.50 \ \text{dollars}}{1 \ \text{bag}} = 3.25 \text{ dollars}\] everything looks good. I agree that if you need 5 pounds of bread, then it will cost you $3.25

OpenStudy (anonymous):

I'm having trouble getting the numbers in the eggs correctly. I have (24 oz) * 1 tub/160 , but I'm not sure what to get as the tub measurement. Could you assist me with that? Please and thanks.

OpenStudy (anonymous):

Would it be 3/20?

jimthompson5910 (jim_thompson5910):

`Eggs, Whole - 1# 8oz` 1 pound, 8 oz = 16 oz + 8 oz = 24 oz so we need 24 oz of whole eggs ------------------------------------------------------- `Eggs, Whole $23.20/tub 10#/tub` 10 pounds = 10*16 = 160 oz 24 oz of whole eggs = (24 oz)*(1 tub/160 oz) = 3/20 tub = 0.15 tub of whole eggs correct

OpenStudy (anonymous):

Would it be 3/20 * 23.20/1 tub = $3.48 for the 24 oz of whole eggs?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Alright, super close! Just a little more. Would you mind assisting me with a smaller question similar to this? I involves the same concept, just phrased differently I believe. I don't think it will take up as much time.

jimthompson5910 (jim_thompson5910):

sure I can help

OpenStudy (anonymous):

I'll get to that one later, the canola oil confuses me a little bit'. What does it mean by 2 ea 17 1/2 ?

OpenStudy (anonymous):

It's the "2 ea" bit that puzzles me.

jimthompson5910 (jim_thompson5910):

"ea" seems like "each" but let me do a bit of thinking and research for a moment

OpenStudy (anonymous):

Maybe it means it's a bundle of 2 containers that both have 280 oz each?

OpenStudy (anonymous):

So 560 oz altogether?

OpenStudy (anonymous):

Yes, because cs I believe converts to "containers" in culinary phrasing.

jimthompson5910 (jim_thompson5910):

I'm guessing "cs" refers to "cups" ? oh it's containers? hmm strange how they use the full word but then abbreviate it

OpenStudy (anonymous):

Looking over it again, you're probably right. I'm just taking educated guesses. Although if it were cups I wonder why they wouldn't just use C?

jimthompson5910 (jim_thompson5910):

I'm reading here http://www.bettycrocker.com/how-to/tipslibrary/charts-timetables-measuring/equivalent-measures-abbreviations that "c" is used for the abbreviation for "cup" so "cs" may be the plural of that

OpenStudy (anonymous):

How odd, I've never seen it like that before. My teacher is pretty forgiving so if I get tiny details incorrect I think he'd understand. Again, this class isn't very math based... We're literally cooking food 99% of the time so I'll do what I can even if things are a little unclear. I still think it comes as a bundle so they'd be 560 oz together. What do you think?

OpenStudy (anonymous):

although when I think about this, I'm not really sure how to write it out.

jimthompson5910 (jim_thompson5910):

I think I found it http://www.beyond.fr/food/convert.html https://answers.yahoo.com/question/index?qid=20100501195226AA4OTFL they're saying that `c.s.` is french for `cuiller à soupe` which is just a fancy way of saying "tablespoon"

jimthompson5910 (jim_thompson5910):

`$54.00/cs` $54 a table spoon sounds like a lot though...hmm

OpenStudy (anonymous):

54 dollars for two containers of canola oil sounds like a lot too, lol.

OpenStudy (anonymous):

2 ea 17 1/2 containers/cs cs I know for a fact also means containers, I'm wondering if he could've misprinted or something. Let's just make an educated guess, this is moreof a shown-work completion assignment so he wont be too particular.

jimthompson5910 (jim_thompson5910):

ok containers makes more sense because the price is way too high if it's $54 per tablespoon

OpenStudy (anonymous):

this canola oil better contain diamonds

jimthompson5910 (jim_thompson5910):

good point but this puzzles me when it says `1/2 # container/cs` a container per container??

OpenStudy (anonymous):

Okay so would it be set like 1 oz * 2 ea 17 1/2 containers/560 ?

jimthompson5910 (jim_thompson5910):

I'm guessing your teacher meant to say "17 1/2 pounds per container" ?

OpenStudy (anonymous):

Let's go with that, it makes more sense than containers per containers.

jimthompson5910 (jim_thompson5910):

ok so if my assumption is correct we need 1 oz of oil and we have this info about the cost of oil $54 per container 17 1/2 pounds per container and idk what to do with the "2 ea" part

OpenStudy (anonymous):

I think that just means we bought two containers, like they came in a bundle maybe?

jimthompson5910 (jim_thompson5910):

oh ok, gotcha

OpenStudy (anonymous):

I think it basically means 1 container is 17 1/2, but since we bought 2 containers the amount would go up by another 17 1/2, which means we would have 560 oz overall.

OpenStudy (anonymous):

But both together they're $54

jimthompson5910 (jim_thompson5910):

1 container = 17 1/2 lbs 2 containers = 35 lbs so you can buy 35 lbs of canola oil for $54 ? That's my guess anyway

OpenStudy (anonymous):

Yes, I think so too

OpenStudy (anonymous):

(1 oz) * 2 containers/560 ?

OpenStudy (anonymous):

or oops, switch those around

OpenStudy (anonymous):

(1 oz) * 560 / 2 containers

OpenStudy (anonymous):

or wait, no I was right the first time.

OpenStudy (anonymous):

so (1 oz) * 2 containers / 560 = 3/1000 ?

OpenStudy (anonymous):

3/1000 container * $54.00/2 containers = .81 cents?

OpenStudy (anonymous):

So altogether, one serving of these stupid egg noodles would be $7.73

jimthompson5910 (jim_thompson5910):

35 lbs = $54 35 lbs = 35*16 = 560 oz 1 oz = (1 oz)*(54 dollars/560 oz) = 0.09642857142858 dollars = 0.10 dollars so 1 ounce of canola oil costs $0.10 or 10 cents that seems kinda low to me, hmm, but then again oil is pretty cheap this page says it's roughly 3.9 cents per fl ounce, so that sounds pretty cheap http://www.walmart.com/ip/Great-Value-Canola-Oil-1-gal/13281828

OpenStudy (anonymous):

I'd believe you're more right, and you don't use too much oil when cooking otherwise things could get yucky.

OpenStudy (anonymous):

So it'd be $7.02, let's stick with that and take a breather.

jimthompson5910 (jim_thompson5910):

tbh if it were me, I'd just buy over the threshold and not worry about the finer details. Besides I don't think you can buy half a container anyway that sounds right though

jimthompson5910 (jim_thompson5910):

what was your other question?

OpenStudy (anonymous):

This one should be more easy, just let me know if you have any idea. I don't want to take up too much of your time since you've literally stuck around for an hour or so. I really appreciate your help, I understand things a lot better now. 5. You take 3 ounces of egg noodle dough and add it to $3.12 worth of tomato sauce and meat to make a spaghetti dinner. If you sell that spaghetti dinner for $9.99 what would be your cost of food sold for the spaghetti dinner?

jimthompson5910 (jim_thompson5910):

I'm guessing all that work beforehand was just to make the egg noodle dough?

OpenStudy (anonymous):

I'm assuming so.

OpenStudy (anonymous):

well, the noodles anyways. I'm not sure if it makes a difference.

jimthompson5910 (jim_thompson5910):

the recipe makes `7# 11 oz` of egg noodles 7# 11 oz = 7*16 + 11 = 123 oz we only need 3 lbs or 3*16 = 48 oz of the noodles full recipe costs $7.02 1 recipe = 7.02 dollars 1 recipe = 123 oz 48 oz = (48 oz)*(1 recipe/123 oz) = 48/123 = 0.39024 recipe 0.39024 recipe = (0.39024 recipe)*(7.02 dollars/1 recipe) = 2.7394848 2.7394848 rounds to 2.74 so for those 48 oz (3 lbs) of noodles used, it costs roughly $2.74

OpenStudy (anonymous):

Added together, they're $5.86. Subtract from 9.99 our profit would be 4.13, correct?

jimthompson5910 (jim_thompson5910):

`You take 3 ounces of egg noodle dough and add it to $3.12 worth of tomato sauce and meat to make a spaghetti dinner` 3 oz of egg noodle = $2.74 tomato sauce + meat = $3.12 total cost per meal = 2.74+3.12 = $5.86

jimthompson5910 (jim_thompson5910):

looks good

OpenStudy (anonymous):

Ah wait, cost. My brain is melting.

OpenStudy (anonymous):

I thought of profit for some reason, ignore that previous statement.

OpenStudy (anonymous):

Alright, that's all.

OpenStudy (anonymous):

Thank you thank you so much

jimthompson5910 (jim_thompson5910):

well I'm sure profit will be asked at some point. A lot of companies obsess about the bottom line

OpenStudy (anonymous):

You've probably taught me better than any math teacher I've ever had, lol.

jimthompson5910 (jim_thompson5910):

I'm glad I could help out

OpenStudy (anonymous):

I really appreciate you taking the time to help me, you're really great. Thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome